OpenStudy (anonymous):

Choose an equation in slope-intercept form for the line that passes through (-12,3) and is perpendicular to the y= -(2/3)x-11.

OpenStudy (anonymous):

a. y = (3/2)x + 3 b. y = (2/3)x - 11 c. y = (2/3)x - 17 d. y = (3/2)x + 21

OpenStudy (anonymous):

first find the slope of the new line. then put the point you have into point-slope form y-y1=m(x-x1)